A marine biologist measures the body length of a bottlenose dolphin to be 343 cm. If the accepted value for the length of the said marine animal is 345 cm, what is the percentage error of the marine biologist/s measurement
1
Expert's answer
2021-09-28T01:03:18-0400

= 345-343 = 2

2/345×100 = 0.6%